Which of the following bonds is different than the others? O (-18,3,3,3, 21) (-45, 7.5, 7.5, 7.5, 50) O (-90, 15, 15, 15, 105please explain your answer
Answer #1

B.(-45,7.5,7.5,7.5,50).

explanation.

consider the following (fifth item / fourth item.)

For option A =21/3 =>7.

for option B =50/7.5 =>6.67.

for option C =105/15=>7

for option D =7/1=>7.

Since for option B it is 6.67 and for others it is 7.

So bond is option B is different than others.